# Distributed Thompson sampling under constrained communication

### v1

@@ Abstract
In Bayesian optimization, a black-box function is maximized via the use of a surrogate model. We apply distributed Thompson sampling, using a Gaussian process as a surrogate model, to approach the multi-agent Bayesian optimization problem. In our distributed Thompson sampling implementation, each agent receives sampled points from neighbors, where the communication network is encoded in a graph; each agent utilizes a Gaussian process to model the objective function. We demonstrate a theoretical bound on Bayesian Simple Regret, where the bound depends on the size of the largest complete subgraph of the communication graph. Unlike in batch Bayesian optimization, this bound is applicable in cases where the communication graph amongst agents is constrained. When compared to sequential Thompson sampling, our bound guarantees faster convergence with respect to time as long as there is a fully connected subgraph of at least two agents. We confirm the efficacy of our algorithm with numerical simulations on traditional optimization test functions, illustrating the significance of graph connectivity on improving regret convergence.
A Introduction
Bayesian optimization (BO) [1, 2] is a technique used to optimize a black-box objective function, where the function is unknown and expensive to evaluate. BO is particularly useful in applications where sampling the objective function involves significant computational resources, such as robotics [3], [4], hyperparameter tuning in machine learning [5], [6], and simulation-based optimization [7]. In BO, a surrogate model is used to approximate the unknown objective function - a Gaussian process (GP) [8] is commonly used for this task. An acquisition function based on this surrogate model is used to determine which points should be queried next. Using the acquisition function, query points can be chosen to incorporate a tradeoff between exploration and exploitation by balancing uncertainty reduction in the model and sampling near the current optimum.
Acquisition functions govern the query selection process, which is critical for ensuring efficient optimization. A popular acquisition function is expected improvement (EI), within which the next query point is selected as one which maximizes the expected improvement over the current best-known value [2]. Alternatively, an upper confidence bound (UCB) can be used to select the next query point by selecting a point which maximizes an upper bound on the predicted value [9]. In entropy search (ES), the query points are selected such that the mutual information between the query points and the location of the global optimum is maximized [10], [11]. In Thompson sampling, the next query point is selected as one which maximizes a realization sampled from the posterior of the surrogate model [12]. Thompson sampling is of particular interest to us, as it lends itself well to multi-agent settings [13].
Evaluation of the efficacy of an implementation of Bayesian optimization is multifaceted. Standard performance metrics for Bayesian optimization include time complexity, sample complexity, and regret. Time complexity describes the number of iterations the algorithm requires to converge to a sufficiently optimal solution [14]. Sample complexity refers to the quantity of samples required for the algorithm to reach an approximately optimal solution [15]. To evaluate the quality of the selection of query points, regret is often used, which quantifies the gap between sampled points and the global optimum [16]. Regret can take many forms, including simple regret, which considers the performance of the best queried point [17], and cumulative regret, which considers performance of all prior sampled points [18].
@@ A Introduction
Bayesian optimization can be used in a multi-agent setting, in which multiple agents can sample the objective function at a single timestep. This is advantageous because in multi-agent settings, the extrema of the objective function can be determined in fewer timesteps than in a single-agent setting - however, additional complexity is introduced, because in order to fully realize this advantage, the agents must engage in collaboration. To facilitate collaboration, existing literature has been focused on batch Bayesian optimization, a centralized setting in which a central coordinator has access to each agents’ acquired information [19], [13]. It then processes and computes the sampling decisions for all agents, which are in turn communicated to each agent. These decisions are often disseminated in batches, allowing multiple agents to simultaneously sample points and consequently parallelizing the optimization process [20], [21].
@@
Centralized approaches are inapplicable in distributed cases, in which there is no centralized coordinator and each agent must possess a local instance of the optimization algorithm [22]. Additionally, they often do not scale well, as they require a central coordinator to manage the processing of all agents’ data. Distributed networks are prevalent in real-world applications, such as in multi-robot source seeking and sensor networks [19], [23]. It may not be the case that all agents have access to all prior sampled points as in the batch setting - communication may be constrained, where some agents are only able to communicate with specific other agents [24]. These constraints may be due to limited communication capacity or computational capacity of the agents, or due to physical proximity constraints. In this work, we study the distributed multi-agent case, where the communication network of agents is given by a static graph.
Our contribution: we propose a distributed Thompson sampling algorithm to approach the multi-agent Bayesian optimization problem. Unlike in batch cases, in which all agents query points based on a posterior distribution which depends on all previously sampled points, our algorithm is applicable in cases with constrained communication. Prior literature providing theoretical guarantees on distributed Thompson sampling algorithms require fully connected communication graphs, even in asynchronous cases [13], [25], and thus are inapplicable in settings with constrained communication. We establish a theoretical bound on regret that is dependent on the size of the largest complete subgraph of the communication graph - we are the first to demonstrate a distributed BO algorithm with probable guarantees in the constrained communication case. Additionally, we establish the efficacy of our performance and confirm the feasibility of our theoretical bound with numerical simulations on standard optimization test functions.
B Problem Formulation and Preliminaries
B-A Problem Formulation
@@
For a compact set $\mathcal{X}\subset\mathbb{R}^{d}$, consider an unknown continuous function $f:\mathcal{X}\to\mathbb{R},$ with optimizer $x^{*}$. In the BO setting, we wish to find the maximum of this function, where we are only able to sample $f$ through expensive and noisy evaluations. We assume any of $M$ agents can query $f$ at any point and receive a noisy value $y=f(x)+\epsilon,$ with $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}).$ Agents query $f$ throughout a total of $T$ iterations. For agent $i\in\{1,\ldots,M\}$ and iteration $t\in\{1,\ldots,T\},$ $x_{t,i}$ is the query point, and $y_{t,i}$ is the corresponding evaluation. Define $X_{t,i}=\{x_{1,i},\ldots,x_{t,i}\},$ $Y_{t,i}=\{y_{1,i},\ldots,y_{t,i}\}$ to be the queries and evaluations made by agent $i$ up to time $t.$ The communication network of $M$ agents is described by graph $G=(V,E),$ where $|V|=M,$ and $E\subset\{\{i,j\}:i,j\in V,i\neq j\}.$ An unordered pair $\{i,j\}\in E$ if agents $i$ and $j$ are able to communicate with each other. Additionally, we denote the set of neighbors of agent $i$ as $N(i)=\{j:\{i,j\}\in E\}.$ The data accessible to agent $i$ at time $t$ is $D_{t,i}=\{(x_{\tau,j},y_{\tau,j})\}_{j\in N(i)\cup i,\tau<t}.$ The set $D_{t,i}$ contains all sampled points by agent $i$ and its neighbors.
B-B Gaussian Process
@@
We will use a Gaussian process (GP) to model our unknown objective function $f$ in our BO setting. Recall the unknown continuous objective function $f:\mathcal{X}\to\mathbb{R}.$ Let $\mathbf{X}_{D_{t}}=\{x_{1},x_{2},\cdots,x_{t}\},$ where $x_{j}$ is the $j$th evaluated point. Let $k:\mathcal{X}^{2}\to\mathbb{R}$ be a kernel function selected based on prior beliefs about $f$. Define
@@
where $\mathbf{K}_{D_{t}}:=[k(x^{\prime},x^{\prime\prime})]_{x^{\prime},x^{\prime\prime}\in\mathbf{X}_{D_{t}}},$ $\mathbf{k}_{D_{t}}(x):=[k(x^{\prime},x)]_{x^{\prime}\in\mathbf{X}_{D_{t}}}$ and $\mathbf{y}_{D_{t}}=\{f(x^{\prime})+\epsilon^{\prime}\}_{x^{\prime}\in\mathbf{X}_{D_{t}}},$ where $\epsilon^{\prime}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}).$ Thus we can define our GP, in which we denote $f|\mathcal{F}_{D_{t}}\sim GP(\mu_{D_{t}}(x),k_{D_{t}}(x,x^{\prime})).$ Note that due to the nature of the $GP$, it is the case that for any $x\in\mathcal{X},$ $f(x)|\mathcal{F}_{D_{t}}\sim N(\mu_{D_{t}}(x),\sigma_{D_{t}}^{2}(x)),$ where $\sigma_{D_{t}}^{2}(x)=k_{D_{t}}(x,x)$ [8]. Furthermore, recall the distributed multi-agent setting, where each of $M$ agents have access to queried points in set $D_{t,i},$ where $D_{t,i}$ and $D_{t,j}$ may not be equal for distinct agents $i$ and $j.$ In our distributed setting, we equip each agent with a unique GP model of $f$, and thus we write $f_{i}(x)|\mathcal{F}_{t,i}\sim\mathcal{N}(\mu_{D_{t,i}}(x),\sigma_{D_{t,i}}^{2}(x))$
B-C Regret
@@
Our metric for algorithm performance is regret, which is an assessment of the quality of sampled points. Specifically, we consider the simple regret, which is the difference between the optimal value of the true function and the best value achieved amongst the previous queried points. This definition of regret is useful because optimization settings focus on locating the extrema of a function, and the simple regret tracks the smallest gap between the value at a sampled point and the optimal value. To account for randomness of $f$ in our regret expression, we take expectation of simple regret to yield the following expression, which is called Bayesian Simple Regret:
$R_{SB}(t)=\min_{i\in\{1,2,\cdots,M\},\tau\in\{1,2,\cdots,t\}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})]$ (1)
@@
In our theoretical analysis, we provide a bound on Bayesian Simple Regret.
B-D Thompson Sampling
@@
Thompson sampling is an algorithm for sequential decision making that can be utilized in this context for determining the next point of the objective function to query [12]. When using Thompson sampling in our Bayesian optimization framework, a function is sampled from the posterior distribution of the Gaussian process. The maximizer of this function is the next query point at which the black-box objective function is sampled. The Gaussian process is then updated with new information from this sample, and the process repeats for the duration of the experiment.
@@
In sequential Thompson sampling, each subsequent query point is determined based on a model updated on all prior sampled points. Alternative implementations of Thompson sampling include batch Thompson sampling, in which multiple query points are determined as a set at each round, and query the objective function in parallel [13], [21]. Batch Thompson sampling is advantageous in systems capable of parallelizing, e.g. multi-agent systems, because it allows for convergence in fewer number of rounds than sequential Thompson sampling.
@@
In the context of multi-agent systems, Batch Thompson sampling can be thought of as a centralized network, where all agents are simultaneously sampling the objective function at different query points, and all agents have access to the same information. However, this may not be realistic in real-world situations, where communication between agents may be more constrained due to bandwidth limitations, computational constrictions, or privacy concerns. In these cases, agents may only have access to the sampled points by some subset of the other agents, and thus each agent will have may have different models of the objective function across their unique GPs. We propose a Thompson Sampling algorithm for this constrained communication case, and provide theoretical guarantees for the algorithm.
C Algorithm: Distributed Thompson Sampling
@@
In our implementation of distributed Thompson sampling, each of $M$ agents have distinct Gaussian processes $\mathcal{GP}_{i}$ for modeling the objective function. At each time step $t,$ all agents update their GPs with the data history available to them. The agent then queries the objective function at $x_{t,i},$ which is the maximizer of the acquisition function sampled from the posterior GP, $\hat{f}_{t,i}\sim\mathcal{GP}_{t,i}.$ Each agent then communicates its sampled point to its neighbors, receives the points sampled by their neighbors, and updates their data history accordingly. The collection of data received by neighbors of agent $i$ at time $t$ is denoted as $C_{t,i}=\{(x_{t,j},y_{t,j})\}_{j\in N(i)}.$ Our method is shown in Algorithm 1.
@@
C-A Theoretical Result
We analyze the performance of the Distributed Thompson Sampling algorithm on the Bayesian Simple Regret metric. Our regret bound depends on the number of timesteps $T$ and the structure of the agent communication graph G. As in prior work, we utilize notions from information theory in our regret bound [26].
Our regret bound involves the Maximum Information Gain (MIG), which is a constant that captures the difficulty of the objective function. MIG has been shown to be bounded for several kernel functions commonly used with GPs, including Squared Exponential and Matérn kernels, the latter of which was used in our numerical implementation [9].

### v3

@@ Abstract
In Bayesian optimization, a black-box function is maximized via the use of a surrogate model. We apply distributed Thompson sampling, using a Gaussian process as a surrogate model, to approach the multi-agent Bayesian optimization problem. In our distributed Thompson sampling implementation, each agent receives sampled points from neighbors, where the communication network is encoded in a graph; each agent utilizes their own Gaussian process to model the objective function. We demonstrate theoretical bounds on Bayesian average regret and Bayesian simple regret, where the bound depends on the structure of the communication graph. Unlike in batch Bayesian optimization, this bound is applicable in cases where the communication graph amongst agents is constrained. When compared to sequential single-agent Thompson sampling, our bound guarantees faster convergence with respect to time as long as the communication graph is connected. We confirm the efficacy of our algorithm with numerical simulations on traditional optimization test functions, demonstrating the significance of graph connectivity on improving regret convergence.
I Introduction
Black-box stochastic optimization involves solving problems where the objective function is not explicitly known and can only be accessed through noisy evaluations [1]. These challenges frequently arise in domains where the evaluation process is costly and uncertain, such as hyperparameter tuning in machine learning [2], [3], simulation-based optimization [4], and experimental design [5]. A variety of methods have been developed to tackle these problems, including evolutionary algorithms [6], particle swarm optimization [7], and finite-difference methods [8]. Among these, Bayesian optimization (BO) [9, 10], has emerged as a particularly powerful framework. In contrast to the aforementioned black-box stochastic optimization algorithms which tend to be model-free, by leveraging a probabilistic surrogate model, often a Gaussian process (GP) [11], BO not only handles the stochastic nature of the evaluations but also balances exploration and exploitation given an appropriately chosen surrogate-based sampling strategy. This data-efficient approach makes BO especially well-suited for optimizing expensive, noisy black-box functions [12]. Moreover, theoretically, BO is also known to satisfy finite-time convergence guarantees to global optima (which we note comes at the cost of a dependence on a term that depends on the complexity of the kernel used to model the underlying function) [12]. To the best of our knowledge, apart from BO, finite-time convergence rates in stochastic optimization are only available for finite-difference type methods. However, due to the local nature of finite-difference type methods, the corresponding finite-time convergence rates for such methods only guarantee convergence to stationary points [8, 13], in contrast to the convergence to global optima achieved by BO algorithms.
In BO, the generation of new sampling points is based on the current surrogate model. A good sampling strategy should balance exploration and exploitation of the current surrogate, which is key for efficient optimization. Common sampling strategies include acquisition function-based approaches such as expected improvement (EI) [10] and BO-upper confidence bound (UCB) [14]. Another popular sampling strategy is Thompson sampling, where the next query point is selected as the optimizer of a random function realization sampled from the current posterior [15, 16]. To evaluate algorithm performance, regret is studied, which quantifies the gap between the performance of sampled points and the global optimum [17]. Types of regret include simple regret, which measures the gap between the optimal value and the performance of the best queried point [18], and cumulative regret, which measures the sum of the gaps between the optimal value and the performance of each queried point [19], [20].
@@ I Introduction
We are interested in multi-agent BO, where multiple agents can sample the objective function at a single timestep. Much of existing multi-agent BO literature studies batch BO, in which a central coordinator has access to each agent’s acquired information [21], [22]. It then computes the sampling decisions for all agents, and communicates these decisions to each agent. These decisions are disseminated in batches, allowing multiple agents to simultaneously sample points, parallelizing the optimization process [23], [24].
@@
Centralized approaches are inapplicable in distributed cases, in which there is no centralized coordinator and each agent must possess a local instance of the optimization algorithm [25]. Additionally, they often do not scale well, as they require a central coordinator to manage the processing of all agents’ data. Distributed networks are prevalent in real-world applications, such as in multi-robot source seeking and sensor networks [21], [26]. It may not be the case that all agents have access to all prior sampled points as in the batch setting - communication may be constrained, where some agents are only able to communicate with specific other agents [27]. These constraints may be due to limited communication capacity or computational capacity of the agents, or due to physical proximity constraints. Prior literature providing theoretical guarantees for distributed Bayesian optimization require fully connected communication graphs, even in asynchronous cases [22], [28], and thus are inapplicable in settings with constrained communication. In this work, we study the distributed setting with constrained communication, in which at each round, agents send their sampled points to their neighbors and receive points sampled by their neighbors.
Our contribution: We propose a distributed Thompson sampling algorithm for the multi-agent Bayesian optimization problem under constrained communication. In the algorithm, each agent uses their own GP model to pick sampling points via Thompson sampling, and shares the queried points with its neighbors. We provide provable guarantees for the proposed distributed BO. In Theorem 3.1, we establish a Bayesian average regret bound of $\tilde{O}\left(\frac{\sqrt{\theta(G)}}{\sqrt{Mt}}\right)$, where $M$ is the number of agents, $t$ is the number of optimization rounds, and $\theta(G)$ represents the clique cover number of $G$, i.e. $\theta(G)$ is the smallest number $L$ such that the graph $G$ can be decomposed into $L$ disjoint complete subgraphs. This implies then that the average regret bound is smaller for graphs with higher connectivity, which can be decomposed into a few large disjoint complete subgraphs. We also characterize Bayesian simple regret, demonstrating a bound of $\tilde{O}\left(\sqrt{\frac{1}{t|V_{max}|}}\right),$ where $|V_{max}|$ is the size of the largest complete subgraph of the communication network $G$. We note that this convergence speed is $O(\sqrt{|V_{max}|})$ times better than the best known simple regret rate for sequential single-agent BO, which is $\tilde{O}\left(\sqrt{\frac{1}{t}}\right)$ [14]. We numerically test our algorithm on two standard optimization test functions [29] with Erdős-Rényi graphs, demonstrating the efficiency of our algorithm. We find that lower regret is achieved with graphs of higher connectivity, supporting our theoretical results.
II Problem Formulation and Preliminaries
II-A Problem Formulation
@@
For a compact set $\mathcal{X}\subset\mathbb{R}^{d}$, consider an unknown continuous function $f:\mathcal{X}\to\mathbb{R},$ with optimizer $x^{*}$. The goal is to find the maximum of this function, where we are only able to sample $f$ through expensive and noisy evaluations. We assume any of $M$ agents can query $f$ at any point and receive a noisy value $y=f(x)+\epsilon,$ with $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}).$ Agents query $f$ throughout a total of $T$ iterations. For agent $i\in\{1,\ldots,M\}$ and iteration $t\in\{1,\ldots,T\},$ $x_{t,i}$ is the query point, and $y_{t,i}$ is the corresponding evaluation. Define $X_{t,i}=\{x_{1,i},\ldots,x_{t,i}\},$ $Y_{t,i}=\{y_{1,i},\ldots,y_{t,i}\}$ to be the queries and evaluations made by agent $i$ up to time $t.$ The communication network of $M$ agents is described by graph $G=(V,E),$ where $|V|=M,$ and $E\subset\{\{i,j\}:i,j\in V,i\neq j\}.$ An unordered pair $\{i,j\}\in E$ if agents $i$ and $j$ are able to communicate with each other. Additionally, we denote the set of neighbors of agent $i$ as $N(i)=\{j:\{i,j\}\in E\}.$ The data accessible to agent $i$ at time $t$ is $D_{t,i}=\{(x_{\tau,j},y_{\tau,j})\}_{j\in N(i)\cup i,\tau<t}.$ The set $D_{t,i}$ contains all sampled points up to time $t$ by agent $i$ and its neighbors. We do not make any assumptions regarding the structure of the communication network. The graph may even be unconnected. Our analysis will show how the graph structure affects the algorithm’s performance.
II-B Gaussian Process
@@
We use a Gaussian process (GP) to model our unknown objective function $f$ in our BO setting. Recall the unknown continuous objective function $f:\mathcal{X}\to\mathbb{R}.$ Let $\mathbf{X}_{D_{t}}=\{x_{1},x_{2},\ldots,x_{t}\},$ where $x_{j}$ is the $j$th evaluated point, and let $k:\mathcal{X}^{2}\to\mathbb{R}$ be a kernel function. Define
@@
where $\mathbf{K}_{D_{t}}:=[k(x^{\prime},x^{\prime\prime})]_{x^{\prime},x^{\prime\prime}\in\mathbf{X}_{D_{t}}},$ $\mathbf{k}_{D_{t}}(x):=[k(x^{\prime},x)]_{x^{\prime}\in\mathbf{X}_{D_{t}}}$ and $\mathbf{y}_{D_{t}}=\{f(x^{\prime})+\epsilon^{\prime}\}_{x^{\prime}\in\mathbf{X}_{D_{t}}},$ where $\epsilon^{\prime}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}).$ Thus we can define our GP, in which we denote $f|\mathcal{F}_{D_{t}}\sim GP(\mu_{D_{t}}(x),k_{D_{t}}(x,x^{\prime})).$ Note that due to the nature of the $GP$, it is the case that for any $x\in\mathcal{X},$ $f(x)|\mathcal{F}_{D_{t}}\sim N(\mu_{D_{t}}(x),\sigma_{D_{t}}^{2}(x)),$ where $\sigma_{D_{t}}^{2}(x)=k_{D_{t}}(x,x)$ [11]. Furthermore, recall the distributed multi-agent setting, where each of $M$ agents have access to queried points in set $D_{t,i},$ where $D_{t,i}$ and $D_{t,j}$ may not be equal for distinct agents $i$ and $j.$ In our distributed setting, each agent $i$ has a unique GP model of $f$ at time $t$, $\mathcal{GP}_{t,i}$, since the data $D_{t,i}$ available to each agent $i$ is different. Thus we denote $f\mid\mathcal{F}_{D_{t,i}}\sim\mathcal{GP}_{t,i}(\mu_{D_{t,i}}(x),k_{D_{t,i}}(x,x^{\prime}))$. In the GP framework, $x^{*},$ the optimizer of $f,$ is treated as a random variable. As a result, $x^{*}$ has a posterior distribution $x^{*}|D_{t},$ which is the optimal value of the GP $f|D_{t}.$ This structure is leveraged in our algorithm.
The kernel function $k(\cdot,\cdot)$ can be selected to reflect prior beliefs about the objective function $f,$ such as function smoothness [11]. Common selections of kernel functions include Linear, Squared Exponential, and Matérn kernels, the latter of which was used in the numerical implementations of our algorithm. Note that the GP problem structure does not make any assumptions regarding function convexity, and that for common kernels, convexity is not reflected by kernel selection.
II-C Regret
@@
Our metric for algorithm performance is regret, which is an assessment of the quality of sampled points. We consider average regret, which quantifies the difference between the optimal value of the function and the queried value for each sampled point. In average regret, this difference is accumulated across all agents and timesteps, and then averaged by the amount of sampled points. To account for randomness of $f$ in our regret expression, we take expectation of average regret to yield the following expression, which we call Bayesian average regret:
$R_{AB}(t)=\frac{1}{tM}\sum_{\tau=1}^{t}\sum_{i=1}^{M}\mathbb{E}[f(x^{*})-f(x_{\tau,i})]$ (1)
We also consider the simple regret, which is the difference between the optimal value of the function and the best value achieved amongst the previous queried points. This definition of regret is useful because optimization settings focus on locating the extrema of a function, and the simple regret tracks the smallest gap between the value at a sampled point and the optimal value. We take the expectation of simple regret to yield the following expression, which is called Bayesian simple regret:
$R_{SB}(t)=\min_{i\in\{1,2,\ldots,M\},\tau\in\{1,2,\ldots,t\}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})]$ (2)
@@
In our theoretical analysis, we provide bounds on Bayesian average regret and Bayesian simple regret.
II-D Thompson Sampling
@@
Thompson sampling is an algorithm for sequential decision making that can be utilized in this context for determining the next point of the objective function to query [15]. When using Thompson sampling in our Bayesian optimization framework, an acquisition function is sampled from the posterior distribution of the Gaussian process. The maximizer of this function is the next query point at which the black-box objective function is sampled. The Gaussian process is then updated with new information from this sample, and the process repeats for the duration of the experiment.
@@
In sequential single-agent Thompson sampling, each subsequent query point is determined based on a single model updated on all prior sampled points. Alternatively, in batch Thompson sampling, multiple query points are determined as a set at each round, and the objective function is sampled in parallel [22], [24]. Batch Thompson sampling is advantageous in systems capable of parallelizing, e.g. multi-agent systems, because it allows for convergence in fewer number of rounds than sequential single-agent Thompson sampling.
@@
Batch Thompson sampling is centralized, with all agents having access to the same information. However, this may not be realistic in real-world situations, where communication between agents may be constrained due to bandwidth limitations, computational constrictions, or privacy concerns. In these cases, agents may only have access to the sampled points by few other agents, and thus datasets available to distinct agents may differ. We propose a distributed Thompson sampling algorithm for this constrained communication case, and provide theoretical guarantees for the algorithm.
III Algorithm: Distributed Thompson Sampling
@@
In our implementation of distributed Thompson sampling, each of $M$ agents have distinct Gaussian processes $\mathcal{GP}_{i}$ for modeling the objective function. At each time step $t,$ all agents update their GPs with the data history available to them. The agent then queries the objective function at $x_{t,i},$ which is the maximizer of the acquisition function sampled from the posterior GP, $\hat{f}_{t,i}\sim\mathcal{GP}_{t,i}.$ Each agent then communicates its sampled point to its neighbors, receives the points sampled by their neighbors, and updates their data history accordingly. The collection of data received by neighbors of agent $i$ at time $t$ is denoted as $C_{t,i}=\{(x_{t,j},y_{t,j})\}_{j\in N(i)}.$ Our method is shown in Algorithm 1. We stress that while we do assume a synchronous global clock, there exists no centralized coordinator in our algorithm that coordinates the queries of the different agents.
@@
In step 11, we select the next sampling point of the objective function by finding the argmax of a function drawn from the posterior distribution of the GP. In our numerical implementation, we did so using gridsearch, but such an approach is computationally expensive for higher dimensional search spaces. Efficient computation of the argmax for Thompson sampling in high dimensional spaces is an active area of research, and a direction for future work.
III-A Theoretical Result
We analyze the performance of the distributed Thompson sampling algorithm on the Bayesian average regret and Bayesian simple regret metrics. Our regret bound depends on the number of timesteps $T$ and the structure of the agent communication graph G. As in prior work, we utilize notions from information theory in our regret bound [30].
Our regret bound involves the Maximum Information Gain (MIG), which is a constant that captures the complexity of the objective function. MIG has been shown to be bounded for several kernel functions commonly used with GPs, including Squared Exponential and Matérn kernels, the latter of which was used in our numerical implementation [14].